OpenStudy (anonymous):

What is the equation of the line that has a slope of 3 and goes through point (4, 7)?

OpenStudy (mathstudent55):

The slope-intercept form of the equation of a line is: \(y = mx + b\)

OpenStudy (mathstudent55):

You are given the slope, 3. In the equation above, m is the slope. You are also given one point, (4, 7). Use the equation above, and let m = 3, x = 4, and y = 7. Then solve for b.

OpenStudy (mathstudent55):

Finally, rewrite the equation above using 3 for m and the value of b you found for b.

OpenStudy (anonymous):

y-7=3(x-4) is what i got

OpenStudy (anonymous):

is that correct?